OpenStudy (anonymous):

x^(3/2)y^(5/3)(x^(7/3)y^(3/2))

OpenStudy (paxpolaris):

I want to be sure, is this what your question is? \[\Huge x^\frac32 \cdot y^\frac53 \cdot x ^\frac73 \cdot y^\frac32\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the last two terms were in ( )

OpenStudy (paxpolaris):

right, but in multiplication, the order in which you multiply doesn't matter....(commutative property of multiplication) so we can regroup the x's and the y's together, like this: \[\Large = \left( x^\frac32 \cdot x ^\frac73 \right) \left( y^\frac53 \cdot y^\frac32 \right)\]

OpenStudy (anonymous):

that is what i was thinking then just add the fractions

OpenStudy (paxpolaris):

yes

OpenStudy (anonymous):

okay thanks alot